Suppose the simplified consolidated balance sheet shown below is for the entire commercial banking 

system. All figures are in billions of rands. The reserve ratio is 9%.

Table 2: Simplified Consolidated Banking System Balance Sheet

Assets Liabilities

Reserves 

Securities

Loans

R70.5

R219.5

R160

Deposits R450 

Based on the information above: The Money multiplier is_______. The value of excess reserves in the 

commercial banking system is ___ billion. If these excess reserves were loaned out, ceteris paribus, the 

total value of additional loans created would be___ billion. 

A. 11,11; R30; R 333

B. 0.09; R30,5; R160

C. 11,11; R30; R1778

D. 40.5; R30; R5,6


1
Expert's answer
2021-09-12T19:31:22-0400

Reserves = required reserves + excess reserves

required reserve ratio "= 9\% = 0.09"

Required reserves "= 0.09 \times 450 = 40.5"

Excess reserves "= 70.5 - 40.5 = 30"

loans created"=\frac{excess\space reserves}{reserve\space ratio}=\frac{30}{0.09}=333.33"

money multiplier"=\frac{1}{0.09}=11.11"